### v2

# Causal Deciphering and Inpainting in Spatio-Temporal Dynamics via Diffusion Model

Abstract
Spatio-temporal (ST) prediction has garnered a De facto attention in earth sciences, such as meteorological prediction, human mobility perception. However, the scarcity of data coupled with the high expenses involved in sensor deployment results in notable data imbalances. Furthermore, models that are excessively customized and devoid of causal connections further undermine the generalizability and interpretability. To this end, we establish a framework for ST predictions from a causal perspective, termed CaPaint, which targets to identify causal regions in data and endow model with causal reasoning ability in a two-stage process. Going beyond this process, we build on the front door adjustment as the theoretical foundation to specifically address the sub-regions identified as non-causal in the upstream phase. By using a fine-tuned unconditional Diffusion Probabilistic Model (DDPM) as the generative prior, we in-fill the masks defined as environmental parts, offering the possibility of reliable extrapolation for potential data distributions. CaPaint overcomes the high complexity dilemma of optimal ST causal discovery models by reducing the data generation complexity from exponential to quasi-linear levels. Extensive experiments conducted on five real-world ST benchmarks demonstrate that integrating the CaPaint concept allows models to achieve improvements ranging from 3.7$\sim$77.3%. Moreover, compared to traditional mainstream ST augmenters, CaPaint underscores the potential of diffusion models in ST data augmentation, offering a novel paradigm for this field. Our project is available at CaPaint.
1 Introduction
Deep learning methodologies have achieved groundbreaking success across a wide array of spatio-temporal (ST) dynamics systems [28, 84, 44], which include meteorological forecasting [3, 50, 59, 92], wildfire spread modeling [71, 20], intelligent transportation [29, 27, 87], and human mobility systems [27, 90], to name just a few. Traditional ST dynamics approaches, based on first-principles [4, 53], often come with high computational costs. In contrast, ST dynamic analysis methods based on deep learning are not directly reliant on the explicit expression of physical laws but are data-driven [28, 84, 3, 27], relying on training models with large-scale observable datasets [86, 65, 92].
In a parallel vein, numerous efforts aim to incorporate physical laws into deep networks [35, 8, 54, 31, 81], termed Physics-Informed Neural Networks (PINNs), which blend deep learning principles with physics to address challenges in scientific computing, particularly in fluid dynamics. PINNs augment traditional neural network models by including a term in the loss function that accounts for the physical laws governing fluid dynamics, such as the Navier-Stokes equations [11]. This ensures that the network’s predictions are not only consistent with empirical data but also comply with the fundamental principles of fluid dynamics. However, the off-the-shelf PINNs often suffer from limited generalization capabilities, primarily due to their customized loss function designs and the neglect of specific network parameter contexts [70, 16].
To date, the data-driven deep models are still dominant in ST dynamical systems, where the numerical simulation methods and PINNs generally lag behind. The reason may stem from the rise of large models [1, 76, 28] and the high costs associated with collecting ST data from sensors [93, 39], which creates a significant conflict between the increasing size of data-hungry models and the uneven, insufficient data collection. To this end, in the ST domain, there is looming research aimed at enhancing the causality and interpretability of models.
Unfortunately, research into causality within the field of ST dynamics is lagging. Although some work has considered causal design, due to specific domain constraints and architectural design, it can only enhance the tailor-made capabilities of the model for specific tasks [95, 38]. Moreover, causal discovery tools [12, 15] applied to ST systems often confront the “curse of dimensionality” issue during dimension reduction, despite their effectiveness in elucidating causal relationships from statistical data [75, 47]. Furthermore, NuwaDynamics [82] for the first time proposed decomposing causal and non-causal regions in ST sequences and enhancing the robustness and generalizability of downstream model training by generating more potential distribution ST sequences through mixup [100]. CauSTG [106] and CaST [95] address the issue of ST distribution shifts by implicitly modeling the time series embeddings and employing intervention techniques to observe these shifts.
Though promising, CauSTG [106] and CaST [95] focus on modeling graph-related data, they lack an understanding of high-dimensional observational data (Dimension $\rm{D}<256$). NuwaDynamics, on the other hand, explores all environments through backdoor adjustments [51], generating a vast number of sequences, which lead to nearly ${\cal O}({T\times{\cal N}_{E}^{{\cal M}\left({\rm{*}}\right)}})$ training complexity ($T$ represents history time step, ${\cal N}_{E}$ and $\cal{M}(*)$ are the number of the environmental patches and mixup, respectively).
Figure 1: Illustration of the CaPaint overview and advantage across SOTA ST causal model on complexity.
In light of this, we propose a general causal structure plugin, termed CaPaint, designed to decipher causal regions in ST data without adding extra computational cost, while intervening in non-causal areas to boost the model’s generalizability and interpretability. Specifically, our method employs a straightforward approach to causal deciphering, utilizing a vision transformer architecture [33] for self-supervised ST data reconstruction. During reconstruction, we leverage attention scores from the self-attention mechanism [23] to map onto important causal patches, thus endowing the model with interpretability. By ranking the entire set of importance scores, we define those with lower scores as environmental patches, which contribute minimally to the model. Building on this, we perform causal interventions in these environmental areas to aid the model in understanding more latent, complex, and imperceptible distributions, thereby enhancing the overall generalizability of the model (see Figure 1). Concretely, we mask trivial regions and perform generation using DDPM [24, 32] fine-tuned on specific ST data, which can also be interpreted as a ST data inpainting approach.
[…]
  • In this paper, we introduce a novel causal structure plugin, CaPaint, which leverages the concept of frontdoor adjustment from causal theory. CaPaint enables various backbone models to learn from a broader distribution of data while providing enhanced interpretability for the models’ predictions.
  • By integrating diffusion generative models with ST dynamics, CaPaint selectively perturbs non-causal regions while maintaining the integrity of core causal areas. This approach generates valuable and reliable data for scenarios where high-quality data are scarce.
  • We conduct extensive experiments across five diverse and representative datasets from different domains, utilizing seven backbone models to assess the effectiveness of the CaPaint method. The empirical results demonstrate that CaPaint consistently enhances performance on all tested datasets and across all backbone models (3.7$\sim$77.3%).
2 Related work & Technical Background
Spatio-temporal Predictive Learning: Various architectures have achieved significant predictive performance in ST domain, which can primarily be categorized as follows: CNN-based models utilize convolutional layers to effectively capture spatial features [45, 48, 77, 9]. RNN-based models, are capable of processing temporal sequence data and are well-suited for understanding temporal changes, showing excellent performance in the prediction of action continuity [69, 80, 86, 72]. GNN-based models effectively capture spatial dependencies and temporal dynamics in data, making them suitable for complex tasks involving geographic locations and temporal changes [44, 25, 36, 17, 99, 98, 83, 14]. Transformer-based models employ self-attention mechanisms to process sequential data in parallel, enhancing the learning of long-term dependencies, and have been used for ST data prediction in complex scenarios [2, 19, 89, 91, 10, 88].
Causal inference: causal discovery algorithms, originally devised for unstructured random vectors [66, 104], have progressively been adapted for ST data analysis [75, 47]. Within the extensive field of deep learning research, the study of causal inference aims to ensure a more stable and robust learning and reasoning paradigm. Recently, an array of techniques has been developed to delve into the nuances of causal features [60, 61, 43, 97], identifying and eliminating spurious correlations [21, 34, 56].
Generative models especially diffusion-based model has gained significant popularity particularly in image and video generation [24, 64, 62]. Sampling optimization algorithms have been used to accelerate the sampling process of diffusion models, significantly reducing the number of steps while improving efficiency. [67, 41]. Additionally, generative models have also been applied to 3D scene generation and point cloud processing, as demonstrated in [40, 30, 73, 74, 22, 63]
Image Inpainting is a technique used to fill in missing or damaged parts of an image. This field can be broadly categorized into the following types. VAE-based methods: These methods leverage Variational Autoencoders to balance diversity and reconstruction [101, 103, 26]. GAN-based methods: Since the introduction of Generative Adversarial Networks, these methods have been widely used for image inpainting [55, 102, 49]. Diffusion model-based methods: Diffusion models have recently shown outstanding performance in image inpainting [46, 68, 57].
3 Methodology
In this section, we systematically introduce causal structure plugin, CaPaint. Initially, we elucidate the methods employed in the upstream phase to delineate causal and non-causal regions (Sec 3.1). Subsequently, we showcase the theoretical underpinnings supporting the CaPaint (Sec 3.2). Building on this causal theory, we further engage in causal intervention within observational data (Sec 3.3). Lastly, we demonstrate how sampling-enhanced ST observations can benefit the complexity of the model’s on-device deployment (Sec 3.4).
Problem Formulation. In ST settings, We represent ST observations as a sequence $\{X_{t}\}_{t=1}^{T}$, where each observation $X_{t}\in\mathbb{R}^{H\times W\times C_{\text{in}}}$ originates from these sequences. Our objective is to predict the trajectory for the forthcoming $K$ steps, denoted as $\{X_{t+1}\}_{t=T}^{T+K}$, with each future state $X_{t+k}$ mapped within $\mathbb{R}^{H\times W\times C_{\text{out}}}$. Here, $H$ and $W$ indicate the spatial grid dimensions, while $C_{\text{in}}$ and $C_{\text{out}}$ define the input and output dimensionality of the observations, respectively.
Figure 2: The details of CaPaint. (Upper.) The initial phase of discovering causal patches. (Bottom.) The update phase designed to eliminate spurious correlation shifts. Following the upstream training of the ViT, a diffusion model is trained in parallel. Using the identified causal patches as conditions, this generative model then performs inpainting for generating multiple sequences.
3.1 Causal Deciphering
To find the causal (non-causal) patches with no labels, we employ a self-supervised reconstruction approach based on the Vision Transformer (ViT) [13] to identify key regions within ST observations. ViT segments the image into multiple patches and calculates the relationships between them using a self-attention mechanism. Due to no label property, we intentionally omit the use of the [Cls] token in classification task and send data into ViT for encouraging “local-to-global” reconstruction.
Specifically, each ST data $X_{t}$, is divided into $N=HW/p^{2}$ patches, where each patch $x_{t}^{patch}\in\mathbb{R}^{N\times(p^{2}\times C_{\text{in}})}$, with $(H,W)$ being the resolution of the original ST data and $(p,p)$ the resolution of each patch. Subsequently, each patch is mapped to a $\rm{D}$-dimensional token through a learnable linear layer, incorporating position embedding to enhance the model’s sensitivity to positional information. These tokens are then fed into successive $L$ stacked transformer blocks, as described in Equation 1:
$\small L\times\left({\rm{X^{\prime}}}=\underbrace{{\rm{X}}+\rm{MSA}\left(LN\left(X\right)\right)}_{\textrm{Multi-head Attention}}\;\;\Rightarrow\;\;X_{\rm{out}}={\rm{X^{\prime}}}+\underbrace{\rm{MLP}\left(LN\left(X^{\prime}\right)\right)}_{\textrm{Residual Connection}}\right)$ (1)
where LN denotes layer normalization, and MLP represents multi-layer perceptron. The upstream self-supervised reconstruction task enables the model to learn intrinsic property of ST data. Navigating the MSA mechanism [78, 96], each patch $x_{t}^{patch}$ derived from the ST observation $X_{t}$ is transformed into queries $q$, keys $k$, and values $v$, and then calculates the relevance of each patch to others, forming a weighted representation that focuses on the most informative parts. The attention weights $A_{i,j}^{h}$ stored in the attention map $A^{h}$ in each head are computed using the scaled dot-product:
[…]
where ${\psi_{tr}}\in\mathbb{R}^{N\times 3D_{h}}$ are the parameter matrices, $D_{h}$ represents the dimension of each head, $Q$, $K$, and $V$ collectively denote the sets of queries $q$, keys $k$, and values $v$. In our approach, the determination of causal patches, is driven by an analysis of the attention maps $A$. Each row in an attention map is normalized and represents the importance of other patches relative to the current patch $x_{t}^{i}$. However, to ascertain the overall importance of each patch across the entire input, we aggregate the contributions by summing the values along the columns of the $A$. To integrate insights across multiple heads, we sum these measures across all heads and then normalize the resultant vector to derive a comprehensive importance score for each patch:
$\small S\in\mathbb{R}^{N}=\text{Softmax}\left(\sum_{h=1}^{H}\sum_{i=1}^{N}A_{i,j}^{h}\right)$ (3)
where $S$ represents the normalized importance score vector, $A_{i,j}^{h}\in A$ denotes the attention that $x_{t}^{i}$ pays to $x_{t}^{j}$ for each head, $H$ is the number of heads. We sort the importance scores in $S$ and select the patches corresponding to the lowest $K$ scores as environmental patches storing in $O_{e}$. The remaining patches are considered causal patches $O_{c}$:
$O_{c}={\rm{Topk}}\left(\lceil\mathcal{C}(S)\times\epsilon\%\rceil,\;\;{\underset{S_{i}\in S}{\arg\max}{\left\{{\rm{set}}\left(\Psi\left(X_{t}\right)\right)\right)\}}}\right)$ (4)
where $C(S)$ is the counting function, $\epsilon$ represents the proportion of patches selected as causal, and $\Psi(X_{t})$ denotes the set of patches in the ST observation $X_{t}$. We identify the causal patches by locating the indices with the highest values in $S$ and define the non-causal parts as the environmental parts. Our goal is to perform causal interventions on the environmental parts.
3.2 Backdoor Adjustment v.s Frontdoor Adjustment
To address issues of ST data scarcity and poor transferability, we examine the evaluation process using a Structural Causal Model (SCM) [52], as shown in Fig 3. We represent abstract data variables by nodes, with directed links symbolizing causality. The SCM illustrates the interaction among variables through a graphical definition of causation, demonstrating the interconnected nature of these elements. As depicted in the left part, NuwaDynamics employs the backdoor adjustment to enhance the model’s generalization performance:
  • ${{\mathcal{X}}_{\mathcal{C}}}\leftarrow{\mathcal{X}}\to{{\mathcal{X}}_{\backslash{\mathcal{C}}}}$. The input ${\mathcal{X}}$ consists of two disjoint parts ${{\mathcal{X}}_{\mathcal{C}}}$ (causal part) and ${{\mathcal{X}}_{\backslash{\mathcal{C}}}}$ (environmental or trivial part).
  • ${{\mathcal{X}}_{\mathcal{C}}}\to{\mathcal{Y}}\nleftarrow{{\mathcal{X}}_{\backslash{\mathcal{C}}}}$. Here, ${{\mathcal{X}}_{\mathcal{C}}}$ represents the sole endogenous parent that determines the ground truth $\mathcal{Y}$. However, in practical scenarios, ${{\mathcal{X}}_{\backslash{\mathcal{C}}}}$ is also employed in predicting $\mathcal{Y}$, which leads to the formation of spurious associations.
Figure 3: Different SCM architectures of SOTA and CaPaint.
In general, a model ${{\mathcal{F}}_{\emptyset}}$ trained using Empirical Risk Minimization (ERM) often struggles to generalize to the test data $\mathcal{D}_{te}\sim\mathbb{P}_{te}$. Such distribution shifts are often induced by variations in environmental patches. Hence, addressing the confounding effect caused by the environmental confounder is crucial. Backdoor adjustment techniques are employed to perturb the environmental components, thereby enhancing the model’s potential to observe a broader range of latent distributions by forcibly perturbing the environmental variables ${{\mathcal{X}}_{\backslash{\mathcal{C}}}}$ (referred to as the do-calculus [51] operator). Unfortunately, traversing all environmental variables is quite challenging. Although NuwaDynamics uses Gaussian sampling to mitigate the issue of complexity, controlling Gaussian sampling in temporal sequence operations is particularly difficult. It requires meticulous adjustment of mean and variance to ensure a balance between the number of environmental samples and the training burden. Worse still, by traversing all environments, it likely violates underlying properties, including distribution shift content and nonexistent scenarios [94]. To address this issue, we employ front-door adjustment, as illustrated in the right half of the Fig 3:
  • ${\tilde{\mathcal{X}}_{\mathcal{C}}}\leftarrow{\mathcal{D}}\to{\mathcal{Y}}$. In this structure, ${\mathcal{D}}$ serves as a confounder, creating a misleading path between ${\tilde{\mathcal{X}}_{\mathcal{C}}}$ and ${\mathcal{Y}}$. Here, ${\tilde{\mathcal{X}}_{\mathcal{C}}}$ represents the causal component within $\tilde{\mathcal{X}}$.
  • ${\tilde{\mathcal{X}}_{\mathcal{C}}}\to\tilde{\mathcal{X}}_{\mathcal{C}}^{*}\to{\mathcal{Y}}$. $\tilde{\mathcal{X}}_{\mathcal{C}}^{*}$ acts as the surrogate variable of ${\tilde{\mathcal{X}}_{\mathcal{C}}}$ and completes ${\tilde{\mathcal{X}}_{\mathcal{C}}}$ to align it with the data distribution. Initially, it derives from and encompasses ${\tilde{\mathcal{X}}_{\mathcal{C}}}$. Specifically, it envisions the potential complete observations that should exist when observing the sub-counterpart ${\tilde{\mathcal{X}}_{\mathcal{C}}}$. Additionally, $\tilde{\mathcal{X}}_{\mathcal{C}}^{*}$ adheres to the data distribution and upholds the intrinsic knowledge of graph properties, thus eliminating any link between ${\mathcal{D}}$ and $\tilde{\mathcal{X}}_{\mathcal{C}}^{*}$. Consequently, $\tilde{\mathcal{X}}_{\mathcal{C}}^{*}$ is well-suited to act as the mediator, which in turn influences the model’s predictions ($\to{\mathcal{Y}}$).
In our front-door adjustment framework, we utilize do-calculus on the variable ${\tilde{\mathcal{X}}_{\mathcal{C}}}$ to eliminate the spurious correlations introduced by ${\mathcal{D}}\to{\mathcal{Y}}$. Specifically, we achieve this by summing over potential surrogate observations $\tilde{X}_{\mathcal{C}}^{*}$. This approach allows us to connect two identifiable partial effects: ${\tilde{\mathcal{X}}_{\mathcal{C}}}\to\tilde{\mathcal{X}}_{\mathcal{C}}^{*}$ and $\tilde{\mathcal{X}}_{\mathcal{C}}^{*}\to{\mathcal{Y}}$:
$P\left({{\mathcal{Y}}{\rm{|}}do\left({{{\tilde{\mathcal{X}}}_{\mathcal{C}}}={{\tilde{X}}_{\mathcal{C}}}}\right)}\right)=\mathop{\sum}\limits_{\tilde{X}_{\mathcal{C}}^{*}}P\left({{\mathcal{Y}}{\rm{|}}do\left({\tilde{\mathcal{X}}_{\mathcal{C}}^{*}=\tilde{X}_{\mathcal{C}}^{*}}\right)}\right)P\left({\tilde{\mathcal{X}}_{\mathcal{C}}^{*}=\tilde{X}_{\mathcal{C}}^{*}{\rm{|}}do\left({{{\tilde{\mathcal{X}}}_{\mathcal{C}}}={{\tilde{X}}_{\mathcal{C}}}}\right)}\right)$ (5)
$={\sum\limits_{{\overset{\sim}{X}}_{\mathcal{C}}^{*}}{\sum\limits_{{\overset{\sim}{X}}_{\mathcal{C}}^{{}^{\prime}}}{P\left(\mathcal{Y}\middle|{{\overset{\sim}{\mathcal{X}}}_{\mathcal{C}}^{*}={\overset{\sim}{X}}_{\mathcal{C}}^{*};{\overset{\sim}{\mathcal{X}}}_{\mathcal{C}}={\overset{\sim}{X}}_{\mathcal{C}}^{{}^{\prime}}}\right)}}}P\left({\overset{\sim}{\mathcal{X}}}_{\mathcal{C}}={\overset{\sim}{X}}_{\mathcal{C}}^{{}^{\prime}}\right)P\left({{\overset{\sim}{\mathcal{X}}}_{\mathcal{C}}^{*}={\overset{\sim}{X}}_{\mathcal{C}}^{*}}\middle|{do\left({{\overset{\sim}{\mathcal{X}}}_{\mathcal{C}}={\overset{\sim}{X}}_{\mathcal{C}}}\right)}\right)$
$={\sum\limits_{{\overset{\sim}{X}}_{\mathcal{C}}^{*}}{\sum\limits_{{\overset{\sim}{X}}_{\mathcal{C}}^{{}^{\prime}}}{P\left(\mathcal{Y}\middle|{{\overset{\sim}{\mathcal{X}}}_{\mathcal{C}}^{*}={\overset{\sim}{X}}_{\mathcal{C}}^{*};{\overset{\sim}{\mathcal{X}}}_{\mathcal{C}}={\overset{\sim}{X}}_{\mathcal{C}}^{{}^{\prime}}}\right)}}}P\left({\overset{\sim}{\mathcal{X}}}_{\mathcal{C}}={\overset{\sim}{X}}_{\mathcal{C}}^{{}^{\prime}}\right)P\left({{\overset{\sim}{\mathcal{X}}}_{\mathcal{C}}^{*}={\overset{\sim}{X}}_{\mathcal{C}}^{*}}\middle|{{\overset{\sim}{\mathcal{X}}}_{\mathcal{C}}={\overset{\sim}{X}}_{\mathcal{C}}}\right)$
$P\left({\tilde{\mathcal{X}}_{\mathcal{C}}^{*}{\rm{|}}do\left({{{\tilde{\mathcal{X}}}_{\mathcal{C}}}={{\tilde{X}}_{\mathcal{C}}}}\right)}\right)=P\left({\tilde{\mathcal{X}}_{\mathcal{C}}^{*}{\rm{|}}{{\tilde{\mathcal{X}}}_{\mathcal{C}}}={{\tilde{X}}_{\mathcal{C}}}}\right)$ holds as ${\tilde{\mathcal{X}}_{\mathcal{C}}}$ is the only parent of $\tilde{\mathcal{X}}_{\mathcal{C}}^{*}$. With data pair $({\tilde{\mathcal{X}}_{\mathcal{C}}},\tilde{\mathcal{X}}_{\mathcal{C}}^{*})$, we can feeding the surrogate observations $\tilde{\mathcal{X}}_{\mathcal{C}}^{*}$ into our ST framework, conditional on the ${\tilde{\mathcal{X}}_{\mathcal{C}}}$, to estimate $P\left(\mathcal{Y}\middle|{{\overset{\sim}{\mathcal{X}}}_{\mathcal{C}}^{*}={\overset{\sim}{X}}_{\mathcal{C}}^{*};{\overset{\sim}{\mathcal{X}}}_{\mathcal{C}}={\overset{\sim}{X}}_{\mathcal{C}}^{{}^{\prime}}}\right)$. Compared to previous work NuwaDynamics, CaPaint utilizes causal regions to generate global surrogate variables in a more rational manner, circumventing the cumbersome need to traverse environmental variables inherent in backdoor adjustments. In fact, backdoor adjustments often likely violate underlying properties, leading to the generation of non-existent data distributions. The broader scenarios of CaPaint will be detailed in Appendix C.
3.3 Causal Intervention via Diffusion Inpainting
Building on the principles of causal analysis outlined above, we proceed to perform interventions on the environmental patches using diffusion inpainting, which enables us to manipulate the environmental areas. Initially, given the unique complexities of ST datasets, we fine-tune the diffusion parameters to adapt seamlessly to the domain-specific challenges, which enhances the accuracy of our interventions on environmental patches. Diffusion models learn the distribution of data through a forward noise addition process and a reverse denoising process:
$\small q(X_{t}\mid X_{t-1})=\mathcal{N}(X_{t};\sqrt{1-\beta_{t}}X_{t-1},\beta_{t}I),\quad p_{\theta}(X_{t-1}\mid x_{t})=\mathcal{N}(X_{t-1};\mu_{\theta}(X_{t},t),\Sigma_{\theta}(X_{t},t))$ (6)
where $X_{t}$ represents the data state at time step $t$, undergoing a transformation from its previous state $x_{t-1}$, $\beta_{t}$ controls the variance of the noise added at each step in the forward process, $\mu_{\theta}$ and $\Sigma_{\theta}$ are neural network outputs that approximate the mean and covariance, respectively. The fine-tuning objective of the diffusion process is designed to approximate the data distribution more accurately. Specifically, the training objective for diffusion models, denoted as $\epsilon_{\theta}$, which predicts the noise, is typically defined as a simplified version of the variational bound:
$L_{\text{simple}}=\mathbb{E}_{X_{0},\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\bm{c},t}\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\big{(}X_{t},\bm{c},t\big{)}\|^{2}$ (7)
where $c$ is the condition information. In this paper, we perform inpainting on the environmental patches of ST data. Inspired by [42], we generate a mask image for each ST data where the causal patches are black and the environmental patches are white. By independently sampling the causal and environmental patches and applying the diffusion inpainting process, we are able to generate augmented ST observation data. The detailed algorithmic process is shown in Appendix A.
$X_{t-1}^{cau}=\sqrt{\bar{\alpha}_{t}}X_{0}+(1-\bar{\alpha}_{t})\epsilon,\quad X_{t-1}^{env}=\frac{1}{\sqrt{\alpha_{t}}}\left(X_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(X_{t},t)+\sigma_{t}z\right)$ (8)
$X_{t-1}=m\odot X_{t-1}^{cau}+(1-m)\odot X_{t-1}^{env}$ (9)
where $X^{cau}$ and $X^{env}$ denote causal patches and environmental patches, $m$ is a binary mask matrix, $\alpha_{t}$ represents the scaling factor at each diffusion step, determining the variance retained in the transition from $X_{t-1}$ to $X_{t}$. The cumulative product $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ represents the accumulated scaling effect from the $T=0$ to step $t$. Equation 9 illustrates the merging of environmental patches and causal patches. Finally, the enhanced ST observation data are stored within our temporal sequence repository to bolster the downstream backbone.
3.4 ST Sequence Sampling Modeling
Previous work [82] assumed that the closer the time point is to the present, the greater its influence, and thus used Gaussian sampling to select more ST data closer to the current time point. However, we argue that uniform sampling can better enhance the model’s generalization ability. To enhance computational efficiency while ensuring prediction accuracy, we employ a ST sequence modeling approach that samples at each time point with a fixed probability controlled by the hyperparameter $p$. This method allows us to sample from both original and generated data at each time point, thereby creating a new spatiotemporal sequence. We use two hyperparameters: $p$, which controls the sampling probability, and $r$, which determines the number of generated spatiotemporal sequences, achieving an optimal balance between computational efficiency and prediction accuracy. The specific sampling process can be represented by the following equation:
$X^{\prime}_{t}=\text{Sample}(X_{t},p,r)$ (10)
where $X_{t}$ represents the collection of original and generated data at time point $t$, and $\text{Sample}(X_{t},p)$ denotes the dataset obtained by sampling from $X_{t}$ with probability $p$. The hyperparameter $p$ is directly set as the sampling probability, while $r$ is used to specify the number of generated ST sequences.
4 Experiments
In this section, we will validate the effectiveness of our proposed causal structure plugin, CaPaint. We design four research questions (RQs) to comprehensively evaluate the performance of CaPaint: RQ1: Does CaPaint effectively enhance model performance and applicability? RQ2: How does CaPaint perform in data-scarce scenarios? RQ3: How does the performance of CaPaint compare with other augmentation methods? RQ4: Is CaPaint effective for long-term time step predictions? Through these research questions, we aim to validate the effectiveness and advantages of CaPaint in handling ST data from multiple perspectives.
4.1 Experimental settings
Datasets. We extensively evaluate our proposal using a diverse range of benchmark datasets spanning multiple fields, include FireSys [7], SEVIR [79], Diffusion reaction system (DRS) [6], KTH [58] and TaxiBJ+ [37]. Specifically, FireSys represents fire dynamics, SEVIR covers meteorological events, DRS involves physical control systems, KTH focuses on human motion dynamics, and TaxiBJ+ is a transportation dataset. Detailed information can be found in the Appendix B.
Backbones and Metrics To validate the generalizability of CaPaint, we select multiple model frameworks for our experiments, including the classic model like ConvLSTM [65], PredRNN-V2 [86], Vision Transformer (ViT) [13], MAU [5], the efficiency-focused SimVP [18], and some of the latest models like MmvP [105] and Earthfarsser [92]. Our evaluation metrics include mean absolute error (MAE), mean squared error (MSE), and structural similarity index measure (SSIM). Detailed information can be found in the Appendix D.
Figure 4: Visualization of prediction results for TaxiBJ+ and SEVIR datasets. The left side shows the predicted results of the last 5 frames for TaxiBJ+. The middle presents the results of long-term predictions for SEVIR, displaying the last five frames from step 10 $\rightarrow$ step 20. The right side compares SSIM metrics with and without the incorporation of CaPaint.
4.2 Evaluating the Efficacy of CaPaint (RQ1 & RQ4)
In this section, we conduct extensive experiments to demonstrate the effectiveness of the CaPaint method. For Transformer architectures, we can directly transfer the model parameters trained in upstream tasks, thereby achieving efficient downstream training. For non-Transformer architectures, we focus on transferring the data itself to train the downstream models. The data presented in the Tab 1 show the performance improvements achieved by generating only one single generalized copy for each ST sequence. As shown in the Table 1, we can list the Observations:
Obs 1. +CaPaint consistently leads w/o Capaint settings across all datasets. As shown in Table 1 and the right side of Fig 4, we can easily observe that introducing +CaPaint significantly improves model performance on MAE, MSE and SSIM metrics across all datasets. For example, with the ViT model on TaxiBJ+, MAE drops from 16.59 $\rightarrow$ 14.54, MSE from 11.40 $\rightarrow$ 8.89; On Diffusion Reaction Systems, MAE significantly decreases from 13.59 $\rightarrow$ 7.52, MSE from 6.21 $\rightarrow$ 1.41. This shows CaPaint’s effectiveness in boosting performance in various domains.
Obs 2. +CaPaint enhances model local insights ST scenarios. By analyzing the left side of Figure 4, we clearly see that the +CaPaint effectively reduces the model’s prediction loss. Moreover, it is observed that +CaPaint provides more accurate predictions in finer details, closely aligning with the actual result curves. This demonstrates CaPaint’s capability to enhance prediction accuracy and reliability, ensuring that the forecasts closely mirror real-world outcomes.
Obs 3. +CaPaint remains effective in long-Term ST predictions. By analyzing the middle of Figure 4, we observe that +CaPaint continues to demonstrate its effectiveness in long-term time step predictions for ST tasks. For instance, the details in the SEVIR dataset predictions improve significantly, indicating that CaPaint is still applicable and beneficial in challenging ST tasks.
Table 1: This table showcases the results (five runs) differences between using the CaPaint concept (+CaP) and not using it (Ori) across various datasets. All MAE and MSE values are multiplied by 100. Blue and Red backgrounds indicate the percentage improvement (reduction) in MAE and MSE, respectively.
Backbone ($10\rightarrow 10$) Metric TaxiBJ+ KTH SEVIR DRS FireSys
Ori +CaP Ori +CaP Ori +CaP Ori +CaP Ori +CaP
ViT [13] MAE $16.59$ $14.54$ $32.03$ $29.52$ $18.69$ $17.56$ $13.59$ $7.52$ $17.32$ $15.97$
MSE $11.40$ $8.89$ $36.11$ $32.79$ $9.93$ $9.16$ $6.21$ $1.41$ $23.40$ $21.06$
$\Delta$ $12.4\%\uparrow$ $22.1\%\uparrow$ $7.8\%\uparrow$ $9.2\%\uparrow$ $6.1\%\uparrow$ $7.7\%\uparrow$ $44.7\%\uparrow$ $77.3\%\uparrow$ $7.8\%\uparrow$ $10.1\%\uparrow$
[…]
MSE $9.94$ $7.83$ $16.84$ $14.07$ $6.75$ $6.01$ $4.09$ $2.24$ $23.37$ $21.94$
$\Delta$ $12.5\%\uparrow$ $21.2\%\uparrow$ $12.6\%\uparrow$ $16.4\%\uparrow$ $5.0\%\uparrow$ $10.9\%\uparrow$ $29.1\%\uparrow$ $37.8\%\uparrow$ $5.1\%\uparrow$ $6.1\%\uparrow$
[…]
MSE $14.22$ $12.29$ $27.31$ $22.37$ $8.45$ $7.26$ $4.11$ $2.32$ $26.09$ $24.97$
$\Delta$ $7.1\%\uparrow$ $13.6\%\uparrow$ $10.0\%\uparrow$ $18.1\%\uparrow$ $16.7\%\uparrow$ $14.1\%\uparrow$ $26.8\%\uparrow$ $43.6\%\uparrow$ $6.2\%\uparrow$ $4.3\%\uparrow$
[…]
MSE $16.79$ $14.67$ $27.37$ $25.15$ $12.12$ $11.41$ $0.64$ $0.31$ $28.64$ $26.44$
$\Delta$ $13.4\%\uparrow$ $12.6\%\uparrow$ $12.1\%\uparrow$ $8.1\%\uparrow$ $10.2\%\uparrow$ $5.9\%\uparrow$ $28.3\%\uparrow$ $51.6\%\uparrow$ $9.6\%\uparrow$ $7.6\%\uparrow$
[…]
MSE $9.60$ $7.89$ $21.45$ $19.11$ $8.54$ $7.73$ $4.37$ $4.18$ $24.71$ $23.12$
$\Delta$ $8.0\%\uparrow$ $16.6\%\uparrow$ $11.6\%\uparrow$ $10.9\%\uparrow$ $9.3\%\uparrow$ $9.4\%\uparrow$ $8.9\%\uparrow$ $4.3\%\uparrow$ $11.6\%\uparrow$ $6.5\%\uparrow$
[…]
MSE $20.46$ $16.60$ $30.19$ $27.84$ $15.43$ $14.34$ $5.28$ $4.66$ $30.89$ $28.91$
$\Delta$ $10.0\%\uparrow$ $18.9\%\uparrow$ $5.9\%\uparrow$ $7.8\%\uparrow$ $3.7\%\uparrow$ $7.1\%\uparrow$ $15.8\%\uparrow$ $11.8\%\uparrow$ $9.8\%\uparrow$ $6.4\%\uparrow$
[…]
MSE $10.96$ $8.21$ $16.46$ $13.91$ $6.82$ $6.21$ $9.54$ $5.03$ $23.34$ $22.11$
$\Delta$ $15.4\%\uparrow$ $25.1\%\uparrow$ $11.4\%\uparrow$ $15.3\%\uparrow$ $5.5\%\uparrow$ $8.9\%\uparrow$ $25.9\%\uparrow$ $47.3\%\uparrow$ $8.4\%\uparrow$ $5.3\%\uparrow$
4.3 Performance in Data-Scarce Scenarios (RQ2)
Figure 5: SSIM improvement across different datasets using the Mmvp model
To assess the performance of CaPaint in data-scarce scenarios, we conducted experiments using varying proportions of training data across multiple datasets and backbones. Specifically, we measured the SSIM improvement at different training data proportions, demonstrating the generalizability and robustness of CaPaint.
Obs 1. CaPaint shows consistent improvements across all training data proportions. As shown in Figures 5 and  6, CaPaint consistently improves SSIM across all training data proportions. This indicates that CaPaint is effective regardless of the amount of training data available, reinforcing its versatility and applicability in diverse scenarios.
Obs 2. Significant performance gains in low data scenarios. The results indicate that CaPaint yields substantial performance improvements, especially in low data scenarios. For instance, with only 10% of the training data, the SSIM improvement is most pronounced, highlighting the method’s effectiveness in data-scarce environments. For example, in the TaxiBJ+ dataset with ViT backbone, the SSIM improvement reaches up to more than 50%, showcasing CaPaint’s capability to enhance model performance with limited data.
Figure 6: SSIM Improvement on DRS across various backbones
Obs 3. Diminishing returns with increased training data. While CaPaint consistently enhances performance, the degree of improvement diminishes as the proportion of training data increases. This trend suggests that the primary benefits of CaPaint are most evident when data is scarce, but the method remains beneficial even as more data becomes available.
Obs 4. CaPaint demonstrates superior performance with equivalent data volumes. As illustrated in Fig 4.4, when comparing 25% original plus 25% augmented data with 50% original data, CaPaint achieves lower MAE and MSE. This demonstrates that CaPaint consistently outperforms the original model by effectively using a mix of original and augmented data, which together match the data volume used by the original model alone.
4.4 Performance Comparison (RQ3)
Table 2: Comparison between CaPaint and other data augmentation methods across various datasets.
Datasets Flip Rotate Crop NuWa CaPaint
DRS $2.10_{\pm 0.16}$ $2.11_{\pm 0.19}$ $2.34_{\pm 0.26}$ $2.02_{\pm 0.09}$ $1.57_{\pm 0.14}$
KTH $23.15_{\pm 1.95}$ $23.14_{\pm 1.67}$ $23.11_{\pm 1.83}$ $22.32_{\pm 0.94}$ $20.56_{\pm 1.02}$
SEVIR $15.41_{\pm 1.49}$ $15.45_{\pm 1.32}$ $15.95_{\pm 1.64}$ $15.14_{\pm 1.57}$ $14.63_{\pm 1.89}$
TaxiBJ+ $16.47_{\pm 0.99}$ $16.39_{\pm 1.32}$ $15.94_{\pm 1.45}$ $15.11_{\pm 0.87}$ $12.87_{\pm 0.76}$
FireSys $17.02_{\pm 2.17}$ $17.07_{\pm 1.94}$ $17.15_{\pm 2.45}$ $16.68_{\pm 1.79}$ $15.79_{\pm 1.88}$
Figure 7: Visualizations in both MAE and MSE with Simvp and + CaP at various training data proportions.
In this section, we compare the performance of different data augmentation methods. Tab 2 shows the model performance using various data augmentation methods across multiple datasets, measured by MAE. It can be seen that traditional data augmentation methods, such as flipping, rotation, and cropping, produce results that are either on par with or slightly worse than the original data. Take the FireSys dataset as an example, MAE increased from 17.01 $\rightarrow$ 17.07 after rotation augmentation. This indicates that conventional data augmentation methods may disrupt the intrinsic properties of ST data, thereby negatively impacting model performance.
In contrast, our method CaPaint achieves the best performance across all datasets. For instance, on the TaxiBJ+ dataset, the MAE with CaPaint augmentation is $12.87$, which is significantly better than the MAE of $15.11$ with NuwaDynamics manual mixup augmentation and the MAE of $15.94$ with other traditional augmentation methods such as cropping. These results highlight the advantage of our method in preserving the integrity of ST data properties. CaPaint not only effectively avoids the disruption caused by data augmentation processes on ST data characteristics but also significantly enhances the model’s predictive capability.
5 Conclusion & Future Work
In this study, we advance the exploration of applying front-door adjustment and causality principles to spatio-temporal forecasting tasks through the introduction of CaPaint. Building upon the foundation of upstream self-supervised learning, we identify causal regions as crucial elements for generating comprehensive and potential data distributions. By integrating diffusion generative models, we ensure the generated data’s rationality and generalizability, thereby enhancing the downstream models’ ability to generalize beyond the observed distribution and improving their interpretability. Moving forward, we plan to explore various generative models for the production of arbitrary-channel ST data to enhance the CaPaint robustness.
6 Acknowledgement
This work was supported by National Natural Science Foundation of China (62476224).
Appendix A CaPaint Inpainting Algorithm
Algorithm 1 Causal Intervention with Diffusion Inpainting
1:Input: ST observation data $X$, masked image $X_{mask}$
2:Output: Augmentation ST observation dataset $X_{A}$
3:Initialize $X_{T}\sim\mathcal{N}(0,I)$ where $T$ is the total number of diffusion steps
4:/* Iterate backwards through diffusion steps */
5:for $t=T$ todo
6:     /* Sample Gaussian noise $\epsilon$/
7:     $\epsilon\sim\mathcal{N}(0,I)$
8:     /* Sample causal region */
9:     $X_{t-1}^{cau}=\sqrt{\bar{\alpha}_{t}}X_{0}+(1-\bar{\alpha}_{t})\epsilon$
10:     /* Sample Gaussian noise $\mathcal{N}$/
11:     $z\sim\mathcal{N}(0,I)$
12:     /* Causal Intervention on Environmental Patches */
13:     $X_{t-1}^{env}=\frac{1}{\sqrt{\alpha_{t}}}\left(X_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(X_{t},t)+\sigma_{t}z\right)$
14:     /* Combine causal and environmental patches */
15:     $X_{t-1}=m\odot X_{t-1}^{cau}+(1-m)\odot X_{t-1}^{env}$
16:     $X_{t}\sim\mathcal{N}(\sqrt{1-\beta_{t-1}}X_{t-1},\beta_{t-1}I)$
17:end for
18:return $X_{A}$ as the augmentation dataset
The algorithm for Causal Intervention with Diffusion Inpainting aims to augment ST observation data through a series of diffusion steps that iteratively refine the data by applying causal interventions and combining them with environmental patches. Here is a detailed step-by-step description:
  • Input: The original ST observation data $X$, and a masked image $X_{\text{mask}}$.
  • Output: An augmented ST observation dataset $X_{A}$.
  • The process begins by initializing $X_{T}$, which represents the data at the final diffusion step, to be a sample from a normal distribution centered at zero with identity covariance.
  • The main loop of the algorithm runs backward from the last diffusion step $T$ to the first. In each step:
    1. 1.
      Gaussian noise $\epsilon_{t}$ is sampled to simulate the diffusion process.
    2. 2.
      A causal region $X_{\text{cau}}$ is sampled where the causal effect is calculated as a blend of the original data and the Gaussian noise, emphasizing areas of interest that should retain more original data characteristics.
    3. 3.
      Gaussian noise $N_{t}$ is sampled again, providing variability to the non-causal or environmental regions.
    4. 4.
      The environmental patches $X_{\text{env}}$ are updated using the data from the previous step adjusted by a damping factor and the added noise, simulating environmental changes.
    5. 5.
      The causal and environmental patches are then combined, where the mask $M$ determines the specific locations for the causal and environmental updates in the data, specifying which parts are from the causal region and which are from the environmental region.
    6. 6.
      The data for the next step, $X_{t-1}$, is computed by normalizing the combined updates, preparing it for the next iteration or output if it is the first step.
  • Finally, the algorithm outputs the augmented data set $X_{A}$, which is the result of the iterative causal intervention and environmental blending over the diffusion process.
Appendix B Details of experiments
SSIM stands for Structural Similarity Index Measure, which is a method for measuring the similarity between two images. It compares the structural information of the images, including luminance, contrast, and texture, to determine how similar they are. SSIM is commonly used in image and video processing applications, such as image compression and quality assessment.
PSNR stands for Peak Signal-to-Noise Ratio. It is a measure of video or image quality that compares the original signal to the compressed or transmitted signal. The higher the PSNR value, the better the quality of the compressed or transmitted signal. PSNR is commonly used in video and image compression applications to evaluate the effectiveness of compression algorithms.
MSE (Mean Squared Error) loss is a commonly used loss function in machine learning and deep learning models. This loss function calculates the average of the squared differences between the predicted and actual values.
Datasets. Here we summarize the details (Tab. 1) of the datasets used in this paper:
  • TaxiBJ+: This dataset contains trajectory data obtained from the GPS of taxis in Beijing, divided into two separate channels: inflow and outflow. Additionally, the dataset has been extended from 32×32 to 128×128 by collecting recent trajectory data from Beijing.
  • KTH: This dataset includes 25 individuals performing six different actions: walking, jogging, running, boxing, waving, and clapping. The complexity of human movements arises from the unique variations each individual displays while executing these actions. By examining previous frames, the model can understand the subtleties of human dynamics and predict future extended postural changes.
  • SEVIR: This dataset consists of weather images that have been sampled and aligned using radar and satellite data. It is designed as a foundational resource to support algorithm development in meteorological research.
  • DRS: This dataset describes the diffusion process of nonlinear wave, which satisfies the diffusion equation.
  • FireSys: The FireSys dataset comprises data associated with fire observations, capturing both temporal and spatial trends of fire evolution, which faithfully represent the progression status in a natural setting.
Appendix C Broader Impact
The development and application of the CaPaint framework in spatio-temporal (ST) dynamics bring several positive broader impacts. Understanding these impacts is crucial for responsible AI research and deployment.
1. Data Imputation in Sparse Scenarios: CaPaint excels in sparse data scenarios, effectively filling in missing data. This reduces the need for extensive sensor deployments, significantly lowering the cost associated with sensor installation. By optimizing data coverage and utilization, CaPaint not only enhances resource efficiency but also achieves substantial cost savings.
2. Enhanced Predictive Accuracy and Interpretability: CaPaint can identify and intervene in non-causal regions, improving the predictive accuracy and interpretability in various ST domains such as meteorology, human mobility, and disaster management. This improvement leads to better decision-making processes and resource allocation, ultimately benefiting society by providing more reliable and understandable predictive models.
3. Cost-Effective Solutions: By reducing the complexity of optimal ST causal discovery models, CaPaint offers a cost-effective solution for handling high-dimensional ST data. This makes advanced predictive technologies more accessible across a broader range of applications, particularly in fields with limited computational resources.
4. Promotion of Causal Reasoning in AI: The integration of causal reasoning into ST models encourages the development of AI systems that better mimic human understanding of cause-and-effect relationships. This can lead to more robust AI models capable of generalizing across different scenarios, fostering trust and reliability in AI applications.
5. Innovation in Data Augmentation Techniques: CaPaint introduces novel data augmentation methods using diffusion inpainting, which can inspire further research and innovation in data augmentation and ST prediction. This can lead to the emergence of new techniques, enhancing the robustness and performance of AI models in various domains.
The CaPaint framework represents a significant advancement in the field of ST dynamics, particularly in its ability to address sparse data scenarios, which reduces the need for extensive sensor deployments and lowers associated costs. Additionally, CaPaint enhances predictive accuracy, interpretability, and efficiency, promotes causal reasoning in AI, and introduces innovative data augmentation techniques. Responsible AI research and deployment should leverage these strengths to maximize benefits while minimizing risks.
Appendix D Metrics
In our research, we investigate the performance of our models using Mean Squared Error (MSE), Mean Absolute Error (MAE), and Structural Similarity Index Measure (SSIM). The formulas for evaluating these indicators, converted into decibels (dB) where applicable, are as follows:
Mean Squared Error (MSE)
Mean Squared Error (MSE) measures the average of the squares of the errors, that is, the average squared difference between the estimated values and the actual value. The MSE is given by:
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}(Y_{i}-\hat{Y_{i}})^{2}$ (D.1)
where $Y_{i}$ is the actual value, $\hat{Y_{i}}$ is the predicted value, and $N$ is the number of observations.
Mean Absolute Error (MAE)
Mean Absolute Error (MAE) measures the average magnitude of the errors in a set of predictions, without considering their direction. It is the average over the test sample of the absolute differences between prediction and actual observation where all individual differences have equal weight. The MAE is given by:
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}\left|Y_{i}-\hat{Y_{i}}\right|$ (D.2)
where $Y_{i}$ is the actual value, $\hat{Y_{i}}$ is the predicted value, and $N$ is the number of observations.
Structural Similarity Index Measure (SSIM)
Structural Similarity Index Measure (SSIM) is used for measuring the similarity between two images. The SSIM index is a decimal value between -1 and 1, where 1 is only reachable in the case of two identical sets of data. The SSIM formula can be quite complex due to its consideration of luminance, contrast, and structure comparison functions between the two images:
$\text{SSIM}(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})}$ (D.3)
where $\mu_{x}$, $\mu_{y}$ are the average of $x$ and $y$ respectively, $\sigma_{x}^{2}$, $\sigma_{y}^{2}$ are the variance of $x$ and $y$ respectively, $\sigma_{xy}$ is the covariance of $x$ and $y$, and $C_{1},C_{2}$ are variables to stabilize the division with weak denominator.
Appendix E Limitations
While the implementation of the CaPaint method has demonstrated significant improvements in prediction accuracy and detail preservation in spatio-temporal forecasting tasks, its enhancements are most pronounced in scenarios characterized by data scarcity or uneven data distribution. In contexts where datasets are abundant and exhibit a broad and uniform distribution, the incremental gains offered by CaPaint may not be as substantial. Nevertheless, the method remains effective, providing consistent, albeit smaller, improvements across diverse data environments.
Appendix F An example of ST Inpainting on SEVIR
Figure F.1: Inpainting Example of our proposed CaPaint.
The figure illustrates the process of maintaining causal regions intact while performing inpainting on non-causal (environmental) regions. The approach involves identifying and deciphering the causal regions (left), intervening by applying diffusion inpainting on the environmental patches (middle), and subsequently generating altered ST data copies (right). This method ensures that the intrinsic causal relationships within the data are preserved, while variations are introduced in the environmental context to augment the dataset effectively.
Appendix G Uneven Distribution of Sensors Leading to Data Scarcity in Global Oceanic Observation
Figure G.1: Temporal distributional heterogeneity within the global oceanic observation platforms, which reveals that there are pronounced disparities in the deployment numbers of various types of sensors during different time intervals.
Appendix H Experimental Parameters
In this experiment, we employ different deep learning models and optimize them for training. All experiments are conducted on hardware equipped with 24 NVIDIA GeForce RTX 4090 GPUs. The optimizer used is Adam, and different learning rates (LR) and batch sizes are set for each model. The specific parameter settings are shown in the table below:
Model Learning Rate (LR) Batch Size
CLSTM 0.001 8
MAU 0.001 8
MMVP 0.004 4
PredRNNv2 0.001 8
SimVP 0.004 4
ViT 0.004 4
Earthfarsser 0.001 8
Table H.1: Learning rates and batch sizes for different backbones
These parameter settings are chosen based on the characteristics of each model and preliminary experimental results on the validation set, aiming to optimize the training efficiency and performance of the models. The Adam optimizer is used with a OneCycle learning rate scheduler, where the maximum learning rate is set according to the specified learning rate for each model, and the number of steps per epoch and the total number of epochs are set based on the training data and experimental setup. During the experiments, we ensure that all models are trained under the same hardware conditions to guarantee the comparability and reproducibility of the results.
Appendix I Visualizations on KTH
Figure I.1: Visualizations on KTH dataset showing the last 5 frames
The first row shows the ground truth for a walking individual. The second row, processed by Earthfarser, exhibits noticeable blurring and loss of detail. The third row, enhanced with +CaPaint, demonstrates a marked improvement in capturing fine details such as the shadow of the person and the accuracy of the foot motion, as highlighted in the red boxes.
Appendix J Visualizations on Diffusion Reaction System
Figure J.1: Visualizations on DRS dataset showing 10 frames
The introduction of CaPaint has led to reductions in Mean Squared Error MSE and MAE, while the SSIM has shown improvements. These changes indicate that the CaPaint method effectively enhances model prediction accuracy and image quality. However, due to the high quality of the model predictions, the improvements might not be readily observable to the naked eye. Despite this, the positive effects of CaPaint are clearly evident through quantitative metrics, demonstrating its potential and practicality in enhancing the accuracy of complex dynamic systems predictions.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In this paper, we introduce the spatio-temporal causal concept in the data mining realm, aimed at enhancing the reliability and accuracy of financial spatio-temporal prediction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In this work, we systematically discuss the limitations of our research and outline directions for future work.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include experimental results related to theoretical aspects.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide the code necessary for replicating the studies described in this paper via an anonymous link, and we detail the experimental setup for the replication in the article itself.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: For the datasets disclosed in the article, we have provided information regarding their sources and origins.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: we have specified all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In this paper, we have reported error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In this paper, we provide detailed information about the experimental resources, including GPU configurations used in our studies.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The study presented in this paper conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We have provided the societal impacts of the work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not address issues related to this aspect.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All creators and original owners of the assets used in our paper, such as code, data, and models, have been properly credited. We have explicitly mentioned the licenses and terms of use for each asset and have ensured full compliance with these terms throughout our research.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The research presented in this paper is not concerned with new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve experiments or research related to human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not address potential risks incurred by study participants.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.